OpenStudy (shelby1290):

Do these systems have one solution, coincidental lines or infinite many solutions and why: 3x-y=0 6x-2y=3

Directrix (directrix):

Multiply this equation by 2 : 3x-y=0 Then, compare it to the second equation.

OpenStudy (shelby1290):

ok so 3 times 2 =6

OpenStudy (shelby1290):

and theres 6x in the second equation...do i have to multiply 1y by 2 as well? @Directrix

Directrix (directrix):

3x times 2 = ? -y times 2 = ? 0 times 2 = ?

OpenStudy (shelby1290):

3x times 2=6x -y times 2=-2y 0 times 2= 0

Directrix (directrix):

How does that compare to the second equation? 6x -2y = 0 6x -2y = 3 ------------

OpenStudy (shelby1290):

that they have the same slope and y-int but a different b

OpenStudy (shelby1290):

does that make it no solution?

Directrix (directrix):

b is used for the y-intercept

Directrix (directrix):

6x -2y = 0 6x -2y = 3 ------------ If you subtract one of these from the other, you get 0 = -3. So, on the day that 0 = -3, the equations will have a solution.

Directrix (directrix):

And, the answer is ? @shelby1290

OpenStudy (shelby1290):

why and what do you subtract its one solution

Directrix (directrix):

Addition/Subtraction, Substitution, and Graphing are 3 ways to find a common solution to a system of 2 equations. So, I subtracted the second from the first one (after you multiplied it by 2) and arrived at something that could not be true. The variables subtracted and disappeared, leaving 0 = -3 which cannot be. No common solution was found. Yes, the lines have the same slope and a different y-intercept.

Directrix (directrix):

>>>that they have the same slope and y-int but a different b What you wrote is not correct. The y-intercepts are different.

OpenStudy (shelby1290):

Kk I see. Thank you!!

OpenStudy (shelby1290):

@Directrix just to check, the final answer is no solution?

Directrix (directrix):

Correct, no solution @shelby1290

OpenStudy (shelby1290):

okay

OpenStudy (shelby1290):

do you have time for another one?

OpenStudy (shelby1290):

@Directrix ^

Directrix (directrix):

Close this thread and post the new question in a fresh thread.

OpenStudy (shelby1290):

okay